### v3

@@ Discussion
For the first external validation set, we utilize the RAD-ChestCT dataset, which includes a cohort from another country [20]. The dataset comprises 3,630 non-contrast chest CT volumes, uniformly reconstructed using a single technique. Half are conducted with Siemens CT, and the other half with GE Medical scanners. The number of slices ranges from 46 to 1277, with an average of 511.6 and a mode of 472. All CT volumes have a resolution of 512 × 512 pixels. The pixel spacing in the axial (XY) plane ranges from 0.189 to 0.977 mm, with an average of 0.692 mm, while the slice thickness varies from 0.125 to 5 mm, with an average of 0.706 mm. To ensure consistent evaluation, we apply identical preprocessing methods to this external dataset as used with our CT-RATE cohort. Since the RAD-ChestCT dataset lacks a Mosaic attenuation label, we exclude it from our external evaluation. Furthermore, while the RAD-ChestCT dataset includes a Calcification label, our dataset distinguishes between Arterial wall calcification and Coronary artery wall calcification. In assessing the model accuracies on the external evaluation set, we use the higher classification probabilities between these two labels for the Calcification prediction.
@@
As a 3D vision transformer, we utilize a component of CT-ViT [23], a network designed to generate low-dimensional tokens from CT volumes and reconstruct CT volumes from these tokens. CT-ViT divides CT volumes into patches measuring $20\times 20\times 10$ voxels in the coronal, sagittal, and axial planes, ensuring uniform spacing. These patches are processed by spatial and causal transformers to produce low-dimensional CT tokens, which are then decoded to reconstruct the CT volumes [23]. For CT-CLIP, we use only the encoder part of this network to generate CT tokens. During the training phase of CT-CLIP, the encoded CT tokens are averaged across the axial plane, then flattened and transformed into a 512-dimensional layer. This creates latent representations for the CT volumes, as shown in Figure 1c. This approach reduces GPU memory requirements significantly through the linear transformation layer for latent representations, while still capturing information from every slice due to the axial attention mechanism in the CT-ViT. The improved memory efficiency enables the use of a larger 3D CT encoder.
For latent representations of radiology reports, we employ the CXR-Bert text encoder [50]. This encoder can process up to 512 tokens, which allows the inclusion of both impression and findings sections from reports. For reports with content of fewer than 512 tokens, we use padding with the $<$PAD$>$ token, where padded tokens are masked during the attention mechanism. Each token is represented in 768 dimensions in the CXR-Bert text encoder. These tokens are then summed across the 512-token span and linearly transformed into a compact 512-dimensional layer, encapsulating the essential textual information from the reports. Figure 1c visually represents this encoding process. Additionally, to assess the impact of including the impression and findings sections, we conduct an ablation study on impression only and findings only training in Supplementary Figure 1a. To highlight the impact of scaling the dataset size on model performance, we conduct an ablation study. This experiment involves training the CT-CLIP model with varying sizes of the CT-RATE dataset: 9.8% (matching the size of the RAD-ChestCT dataset, the only public chest CT dataset with multi-abnormality labels [20]), 20%, 40%, 60%, 80%, and 100%. By systematically altering the sizes of the dataset, our goal is to explore how the accuracy of the model scales with the amount of available training data. Figure 4c illustrates the outcomes of this ablation study.
@@
In assessing our CT-CLIP and finetuned models, we employ the CT-Net model as a fully supervised baseline [20]. Renowned for its performance in a similar CT dataset [20], CT-Net provides a robust comparison point. We train the CT-Net model on our CT-RATE dataset, adapting the chest CT volumes to fit the model’s input requirements by resizing them to $420\times 420\times 402$. Consistent with our models, we focus on the abnormalities present in the CT-RATE dataset (Figure 3d), instead of the 98 and 9 abnormalities explored in the original CT-Net study. For each abnormality, we determine classification thresholds by identifying the left-upper point on the Receiver Operating Characteristic (ROC) curve, using the 10% of each validation set. This analysis is performed across all three evaluation sets—CT-RATE (internal), RAD-ChestCT (external), and UPMC (external)—and for all multi-abnormality detection models: zero-shot, fine-tuned, and fully supervised. Additionally, we compute the Area Under the Receiver Operating Characteristic (AUROC) values for each abnormality.
To provide a comprehensive evaluation, we also calculate accuracy, precision, and weighted F1 scores for each identified pathology. These additional metrics offer more insights into the performance of the models. To further validate the latent spaces generated by CT-CLIP for radiology reports and volumes, we conduct a t-SNE (t-distributed stochastic neighbor embedding) analysis [56]. This analysis enables us to map the positions of abnormalities visually, the distribution of abnormality groups, and other relevant attributes within the t-SNE embeddings (Figures 4d, 4e and Extended Data Figures 1, 2).
@@
[…]
Where $n$ is the total number of samples. It is important to note that as $K$ increases, MAP@K may decrease due to the inclusion of less relevant images. We utilize the vision encoders of the supervised baseline, CT-CLIP, and the finetuned models for volume-to-volume retrieval and assess their performance across both internal and external validation sets using MAP@1, MAP@5, MAP@10, and MAP@50 metrics. The Recall@K metric is not applied in image-to-image retrieval due to the inherent limitation that the cosine similarity for identical images would inevitably be the highest, resulting in a consistent Recall@K value of $1$. However, this issue does not arise in text-to-image retrieval since the query is based on text embeddings rather than image embeddings.
For report-to-volume retrieval, we utilize the pretrained text transformer and 3D vision transformer to generate vector embeddings for both radiology reports and CT volumes, which are transformed from the token space to the embedding space with the CT-CLIP’s trained projection layers (Figure 1c). Our goal is to find the CT volume that best matched each query report. To achieve this, we compute the cosine similarity between the latent representations of each report and all the CT volumes. Subsequently, we rank the CT volumes based on their cosine similarity scores to identify and return the most relevant matches. The process of report-to-volume retrieval is depicted in Figure 5b.
For the evaluation of this task, we employ the Recall@K metric, where $K$ represents the number of returned images. This method serves as a standard evaluation tool in text-to-image retrieval tasks [44]. We assess four Recall@K metrics: Recall@5, Recall@10, Recall@50, and Recall@100. These metrics measure the frequency of the target image appearing among the top 5, 10, 50, and 100 retrieved images, respectively. It is worth noting that the external validation set lacks text reports. Therefore, we restrict the report-to-volume retrieval evaluation to the internal validation set. We assess both CT-CLIP and the finetuned model with VocabFine for this task. However, since the supervised model and ClassFine do not generate latent embeddings for reports, they are not utilized in the report-to-volume retrieval.
@@
The validation set is curated similarly to the training set, using the free-form reports from the CT-RATE validation set. The validation dataset includes 33,326 conversation questions, 8,886 description questions, 8,888 free-response questions, 12,013 short-answer questions, 8,915 multiple-choice questions, and 3,039 report-generation questions. Extended Data Figure 7 shows example entries from the described dataset. Since the VQA dataset is generated by a large language model, we carefully select only accurately generated and properly formatted data—that is, data that can be successfully parsed as JSON—for training our models. Additionally, we limit the initial number of generated questions to maintain a reasonable training duration. Our VQA dataset is made publicly available alongside CT-RATE.
@@
In our implementation, we evaluate and compare four different LLMs. For the primary configuration of CT-CHAT, we use the 70B Meta Llama 3.1 [45], one of the state-of-the-art open-source LLMs. This model consists of 80 transformer layers, 64 attention heads, an embedding dimension of 8,192, an FFN dimension of 28,672, and supports a context size of up to 128k tokens. For a smaller model, we utilize the 8B Meta Llama 3.1 [45], which includes 32 transformer layers, 32 attention heads, an embedding dimension of 4,096, an FFN dimension of 14,336, and a context size of up to 128k tokens. Additionally, for comparison, we implement two LLMs used in the official LLaVA 1.6 [26]: Vicuna 13B and Mistral 7B. Figure 6d shows the scores for each of these LLMs. To dynamically adjust the model’s output mode based on the specific task, we add four special tokens to the tokenizers of each model: <long_answer>, <short_answer>, <report_generation>, and <multiple_choice>. These tokens are selected for different tasks during training and inference. The multimodal projector, based on the approach described in [14], consists of an attention-pooling layer followed by a 2-layer multi-layer perceptron (MLP). The attention-pooling mechanism utilizes 256 learned latent queries in conjunction with multi-headed cross-attention to compress the final layer’s features from the encoder backbone into a fixed-length sequence of image tokens. This approach optimizes both training and inference, ensuring that the sequence length remains within the LLM’s context window, and making the model focus more on the important features with the attention mechanism. The MLP that follows is adapted from LLava 1.6 [26] and includes a single hidden layer activated by GeLU, transforming the image tokens to match the LLM’s embedding dimension.
@@
Training CT-CHAT involves two steps. First, in the medical concept alignment step, the LLM is frozen, and the multimodal projector is trained with the input “Please provide the radiology report for the following chest CT volume <image>” and the output report. This step improves subsequent finetuning performance and helps the multimodal projector to be aligned in biomedical concepts [27]. We conduct an ablation study of this step to show its effect in Supplementary Table 10b. In the second step, medical instruction tuning, the model is trained with the CT-VQA dataset that includes long answer, short answer, multiple choice, and report generation question-answer pairs. During this step, the multimodal projector, along with Low-Rank Adaptation (LORA) weights [46] for the LLM with rank 128 and alpha 256, is finetuned. <image> tokens are added to indicate the embedded image input to the model, except for text-only questions and answers. The four special tokens are added to the respective question-answer pairs in the dataset, enabling the model to learn different output modes. These tokens are later used during inference to select different modes of the model. The hyperparameters for each model for pretraining and finetuning are provided in Supplementary Table 4a, 4b.
@@
Given the lack of pretrained VQA models for 3D imaging, we conduct a comparative analysis of CT-CHAT against four 2D state-of-the-art open-source VQA model: LLaVA 1.6 (Mistral 7B) [26], LLaVA 1.6 (Vicuna 13B) [26], LLaVA-Med [27], and CXR-LLaVA [28], using benchmarking framework established in [14]. To address the challenge of representing 3D information using their 2D encoders and to ensure a fair comparison, we utilize Digitally Reconstructed Radiographs (DRRs) derived from CT-RATE, as suggested by [47]. This approach is chosen over the selection of random or central CT slices, which can not adequately capture all relevant pathologies. The use of DRRs is justified, as [47] demonstrates that models trained with DRRs show comparable results to those trained with 3D chest CTs.
@@
In addition to benchmarking CT-CHAT on VQA tasks, we further evaluate its performance in radiology report generation—a task more reflective of routine radiological workflows—by comparing it with two state-of-the-art 3D report generation models: RadFM [29] and CT2Rep [30]. While VQA metrics such as BLEU, ROUGE, and the Llama score help assess linguistic and clinical quality, they may not fully capture diagnostic utility. Therefore, for this comparison, we employ clinical efficacy metrics—precision, recall, and F1 score—using our automated report labeler, which extracts structured abnormality labels from generated reports. Evaluation is conducted on both the internal CT-RATE validation set and the external RAD-ChestCT dataset. For RAD-ChestCT, which lacks text reports, we leverage its binary abnormality annotations to enable indirect evaluation through our labeler. Importantly, we use the original pretrained weights for both RadFM and CT2Rep without retraining.
CT2Rep is trained on the same CT-RATE dataset as CT-CHAT and uses the same architecture for its vision encoder. However, unlike CT-CHAT, which benefits from a contrastively pretrained CT-CLIP encoder, CT2Rep is trained from scratch without any pretraining. Additionally, CT2Rep is trained specifically for report generation, while CT-CHAT is trained more broadly across multiple VQA tasks—including classification, question-answering, and report generation. Despite this, CT-CHAT consistently outperforms CT2Rep across all clinical efficacy metrics. This performance gap highlights the substantial value of contrastive pretraining with CT-CLIP in capturing clinically meaningful representations of 3D CT volumes, which can then be effectively transferred to downstream tasks such as report generation—even when the model is not trained exclusively for that task.
@@
Our CT-RATE dataset, comprising paired chest CT volumes and radiology text reports, is publicly accessible via https://huggingface.co/datasets/ibrahimhamamci/CT-RATE. The external validation set, RAD-ChestCT, is also openly accessible through https://zenodo.org/records/6406114.
@@
Our trained models and codebase are publicly available at https://github.com/ibrahimethemhamamci/CT-CLIP and https://github.com/ibrahimethemhamamci/CT-CHAT for further research.

### v4

@@ Discussion
For the first external validation set, we use the RAD-ChestCT dataset, which includes a cohort from another country [20]. The dataset comprises 3,630 non-contrast chest CT volumes, uniformly reconstructed using a single technique. Half are conducted with Siemens CT, and the other half with GE Medical scanners. The number of slices ranges from 46 to 1277, with an average of 511.6 and a mode of 472. All CT volumes have a resolution of 512 × 512 pixels. The pixel spacing in the axial (XY) plane ranges from 0.189 to 0.977 mm, with an average of 0.692 mm, while the slice thickness varies from 0.125 to 5 mm, with an average of 0.706 mm. To ensure consistent evaluation, we apply identical preprocessing methods to this external dataset as used with our CT-RATE cohort. Since the RAD-ChestCT dataset lacks a Mosaic attenuation label, we exclude it from our external evaluation. Furthermore, while the RAD-ChestCT dataset includes a Calcification label, our dataset distinguishes between Arterial wall calcification and Coronary artery wall calcification. In assessing the model accuracies on the external evaluation set, we use the higher classification probabilities between these two labels for the Calcification prediction.
@@
As a 3D vision transformer, we use a component of CT-ViT [23], a network designed to generate low-dimensional tokens from CT volumes and reconstruct CT volumes from these tokens. CT-ViT divides CT volumes into patches measuring $20\times 20\times 10$ voxels in the coronal, sagittal, and axial planes, ensuring uniform spacing. These patches are processed by spatial and causal transformers to produce low-dimensional CT tokens, which are then decoded to reconstruct the CT volumes [23]. For CT-CLIP, we use only the encoder part of this network to generate CT tokens. During the training phase of CT-CLIP, the encoded CT tokens are averaged across the axial plane, then flattened and transformed into a 512-dimensional layer. This creates latent representations for the CT volumes, as shown in Figure 1c. This approach reduces GPU memory requirements significantly through the linear transformation layer for latent representations, while still capturing information from every slice due to the axial attention mechanism in the CT-ViT. The improved memory efficiency enables the use of a larger 3D CT encoder.
For latent representations of radiology reports, we employ the CXR-Bert text encoder [49]. This encoder can process up to 512 tokens, which allows the inclusion of both impression and findings sections from reports. For reports with content of fewer than 512 tokens, we use padding with the <<PAD>> token, where padded tokens are masked during the attention mechanism. Each token is represented in 768 dimensions in the CXR-Bert text encoder. These tokens are then summed across the 512-token span and linearly transformed into a compact 512-dimensional layer, encapsulating the essential textual information from the reports. Figure 1c visually represents this encoding process. Additionally, to assess the impact of including the impression and findings sections, we conduct an ablation study on impression only and findings only training in Supplementary Figure 1a. To highlight the impact of scaling the dataset size on model performance, we conduct an ablation study. This experiment involves training the CT-CLIP model with varying sizes of the CT-RATE dataset: 9.8% (matching the size of the RAD-ChestCT dataset, the only public chest CT dataset with multi-abnormality labels [20]), 20%, 40%, 60%, 80%, and 100%. By systematically altering the sizes of the dataset, our goal is to explore how the accuracy of the model scales with the amount of available training data. Figure 4c illustrates the outcomes of this ablation study.
@@
In assessing our CT-CLIP and finetuned models, we employ the CT-Net model as a fully supervised baseline [20]. Renowned for its performance in a similar CT dataset [20], CT-Net provides a robust comparison point. We train the CT-Net model on our CT-RATE dataset, adapting the chest CT volumes to fit the model’s input requirements by resizing them to $420\times 420\times 402$. Consistent with our models, we focus on the abnormalities present in the CT-RATE dataset (Figure 3d), instead of the 98 and 9 abnormalities explored in the original CT-Net study. For each abnormality, we determine classification thresholds by identifying the left-upper point on the Receiver Operating Characteristic (ROC) curve, using the 10% of each validation set. This analysis is performed across all three evaluation sets, CT-RATE (internal), RAD-ChestCT (external), and UPMC (external), and for all multi-abnormality detection models: zero-shot, fine-tuned, and fully supervised. Additionally, we compute the Area Under the Receiver Operating Characteristic (AUROC) values for each abnormality.
To provide a comprehensive evaluation, we also calculate accuracy, precision, and weighted F1-scores for each identified abnormality. For the weighted F1, each of the 18 abnormalities is treated as a binary task (present vs. absent); we compute a weighted F1-score that balances positive and negative cases within each abnormality, then average these scores to produce a single overall metric. This macro of weighted F1 scores ensures that (i) each abnormality contributes equally to the final score and (ii) class imbalance within each abnormality is addressed fairly, resulting in a clinically meaningful measure of performance. To further validate the latent spaces generated by CT-CLIP for radiology reports and 3D chest CT volumes, we also conduct a t-SNE (t-distributed stochastic neighbor embedding) analysis [55]. This allows us to visually map the distribution of abnormalities and other relevant attributes within the t-SNE embeddings (Figures 4d,4e and Extended Data Figures12).
@@
$\text{MAP@K}=\frac{1}{n}\text{AP@K}$
Where $n$ is the total number of samples. It is important to note that as $K$ increases, MAP@K may decrease due to the inclusion of less relevant images. We use the vision encoders of the supervised baseline, CT-CLIP, and the finetuned models for volume-to-volume retrieval and assess their performance across both internal and external validation sets using MAP@1, MAP@5, MAP@10, and MAP@50 metrics. The Recall@K metric is not applied in image-to-image retrieval due to the inherent limitation that the cosine similarity for identical images would inevitably be the highest, resulting in a consistent Recall@K value of $1$. However, this issue does not arise in text-to-image retrieval since the query is based on text embeddings rather than image embeddings.
For report-to-volume retrieval, we use the pretrained text transformer and 3D vision transformer to generate vector embeddings for both radiology reports and CT volumes, which are transformed from the token space to the embedding space with the CT-CLIP’s trained projection layers (Figure 1c). Our goal is to find the CT volume that best matched each query report. To achieve this, we compute the cosine similarity between the latent representations of each report and all the CT volumes. Subsequently, we rank the CT volumes based on their cosine similarity scores to identify and return the most relevant matches. The process of report-to-volume retrieval is depicted in Figure 5b.
For the evaluation of this task, we employ the Recall@K metric, where $K$ represents the number of returned images. This method serves as a standard evaluation tool in text-to-image retrieval tasks [43]. We assess four Recall@K metrics: Recall@5, Recall@10, Recall@50, and Recall@100. These metrics measure the frequency of the target image appearing among the top 5, 10, 50, and 100 retrieved images, respectively. It is worth noting that the external validation set lacks text reports. Therefore, we restrict the report-to-volume retrieval evaluation to the internal validation set. We assess both CT-CLIP and the finetuned model with VocabFine for this task. However, since the supervised model and ClassFine do not generate latent embeddings for reports, they are not used in the report-to-volume retrieval.
@@
The validation set is curated similarly to the training set, using the free-form reports from the CT-RATE validation set. The validation dataset includes 33,326 conversation questions, 8,886 description questions, 8,888 free-response questions, 12,013 short-answer questions, 8,915 multiple-choice questions, and 3,039 report-generation questions. Extended Data Figure 7 shows example entries from the described dataset. Since the VQA dataset is generated by a large language model, we carefully select only accurately generated and properly formatted data, that is, data that can be successfully parsed as JSON for training our models. Additionally, we limit the initial number of generated questions to maintain a reasonable training duration. Our VQA dataset is made publicly available alongside CT-RATE.
@@
In our implementation, we evaluate and compare four different LLMs. For the primary configuration of CT-CHAT, we use the 70B Meta Llama 3.1 [44], one of the state-of-the-art open-source LLMs. This model consists of 80 transformer layers, 64 attention heads, an embedding dimension of 8,192, an FFN dimension of 28,672, and supports a context size of up to 128k tokens. For a smaller model, we use the 8B Meta Llama 3.1 [44], which includes 32 transformer layers, 32 attention heads, an embedding dimension of 4,096, an FFN dimension of 14,336, and a context size of up to 128k tokens. Additionally, for comparison, we implement two LLMs used in the official LLaVA 1.6 [26]: Vicuna 13B and Mistral 7B. Figure 6d shows the scores for each of these LLMs. To dynamically adjust the model’s output mode based on the specific task, we add four special tokens to the tokenizers of each model: <long_answer>, <short_answer>, <report_generation>, and <multiple_choice>. These tokens are selected for different tasks during training and inference. The multimodal projector, based on the approach described in [14], consists of an attention-pooling layer followed by a 2-layer multi-layer perceptron (MLP). The attention-pooling mechanism uses 256 learned latent queries in conjunction with multi-headed cross-attention to compress the final layer’s features from the encoder backbone into a fixed-length sequence of image tokens. This approach optimizes both training and inference, ensuring that the sequence length remains within the LLM’s context window, and making the model focus more on the important features with the attention mechanism. The MLP that follows is adapted from LLava 1.6 [26] and includes a single hidden layer activated by GeLU, transforming the image tokens to match the LLM’s embedding dimension.
@@
Training CT-CHAT involves two steps. First, in the medical concept alignment step, the LLM is frozen, and the multimodal projector is trained with the input “Please provide the radiology report for the following chest CT volume <image>” and the output report. This step improves subsequent finetuning performance and helps the multimodal projector to be aligned in biomedical concepts [27]. We conduct an ablation study of this step to show its effect in Supplementary Table 10b. In the second step, medical instruction tuning, the model is trained with the CT-VQA dataset that includes long answer, short answer, multiple choice, and report generation question-answer pairs. During this step, the multimodal projector, along with Low-Rank Adaptation (LORA) weights [45] for the LLM with rank 128 and alpha 256, is finetuned. <image> tokens are added to indicate the embedded image input to the model, except for text-only questions and answers. The four special tokens are added to the respective question-answer pairs in the dataset, enabling the model to learn different output modes. These tokens are later used during inference to select different modes of the model. The hyperparameters for each model for pretraining and finetuning are provided in Supplementary Table 4a, 4b. To boost report-generation accuracy, we finetune CT-CHAT with structured metadata extracted exclusively from pulmonary-nodule masks. Each CT-RATE volume is processed with TotalSegmentator [61], from which we record, per nodule, the lobar location, centroid, orthogonal diameters, and volume (mm3). These fields are embedded in a fixed prompt:
<volume_tokens>
Generate a detailed radiology report for the given chest CT volume.
<BEGIN_NODULE_INFO>
Nodule_Count: n
Nodule_ID: 1  Lobe: left lower
 Centroid_mm: (x,y,z)
 Diameter_mm: (x:dx, y:dy, z:dz)
 Volume_mm3: v
… (repeat for each nodule)
<END_NODULE_INFO>
Each sample includes (i) CT volume features from the frozen CT-CLIP encoder, (ii) the structured nodule template, and (iii) the reference report. Finetuning is performed on the 8B Llama 3.1 configuration of CT-CHAT using LoRA (rank 128, $\alpha=256$); only the LoRA weights and projector are updated. Supplying these segmentation-derived cues markedly improves performance (Figure 6e). The same workflow can incorporate additional abnormalities as more 3D segmentation models become available.
@@
Given the lack of pretrained VQA models for 3D imaging, we conduct a comparative analysis of CT-CHAT against four 2D state-of-the-art open-source VQA model: LLaVA 1.6 (Mistral 7B) [26], LLaVA 1.6 (Vicuna 13B) [26], LLaVA-Med [27], and CXR-LLaVA [28], using benchmarking framework established in [14]. To address the challenge of representing 3D information using their 2D encoders and to ensure a fair comparison, we use Digitally Reconstructed Radiographs (DRRs) derived from CT-RATE, as suggested by [46]. This approach is chosen over the selection of random or central CT slices, which can not adequately capture all relevant pathologies. The use of DRRs is justified, as [46] demonstrates that models trained with DRRs show comparable results to those trained with 3D chest CTs.
@@
In addition to benchmarking CT-CHAT on VQA tasks, we further evaluate its performance in radiology report generation, a task more reflective of routine radiological workflows, by comparing it with two state-of-the-art 3D report-generation models: RadFM [29] and CT2Rep [30]. While VQA metrics such as BLEU, ROUGE, and the Llama score assess linguistic and clinical quality, they do not fully capture diagnostic use. Therefore, for this comparison we employ clinical-efficacy metrics: precision, recall, F1, and the CRG score [66]. CRG weights findings by prevalence and severity, giving a distribution-aware score aligned with radiologists’ priorities. Metric computation uses our automated report labeler, which extracts structured abnormality labels from the generated reports. Evaluation is conducted on both the internal CT-RATE validation set and the external RAD-ChestCT dataset; for RAD-ChestCT, which lacks text reports, we leverage its binary abnormality annotations to enable indirect evaluation through our labeler. Notably, we use the original pretrained weights for both RadFM and CT2Rep without retraining.
To complement automatic metrics, we performed a blinded, three-reader study of report quality. Two board-certified thoracic radiologists from Zurich (R1, R2) and one from Istanbul (R3) independently rated reports generated by CT-CHAT, RadFM, and CT2Rep. Because free-text scoring is time-intensive, we limited the pool to 150 CT-RATE validation cases per model. Cases were stratified rather than randomly sampled: after computing Llama scores for CT-CHAT outputs, we selected 50 volumes from each accuracy band (low 0–3, medium 4–6, high 7–10) to ensure balanced difficulty. For every selected scan, the three anonymized model outputs were presented, and readers scored each report on a 0–10 scale against predefined criteria covering finding description, localisation, terminology, and overall clinical adequacy. Figure 6d shows the results: CT-CHAT achieved mean scores of 3.44 (R1), 4.10 (R2), and 4.52 (R3), outperforming RadFM (0.55, 0.91, 0.86) and CT2Rep (2.66, 2.52, 2.74). Violin plots display a right-shifted distribution for CT-CHAT, confirming superior perceived clinical fidelity. The same protocol can be repeated for larger or external cohorts as new models emerge. It is important to note CT-CHAT’s superiority over CT2Rep, even though both models are trained on the same CT-RATE dataset. Moreover, CT2Rep is optimized solely for report generation, while CT-CHAT is tuned primarily for VQA tasks, classification, question-answering, and only secondarily for report generation. Despite this, CT-CHAT consistently outperforms CT2Rep across all evaluations. The performance gap underscores the value of contrastive pretraining with CT-CLIP in capturing clinically meaningful 3D CT representations that transfer effectively to downstream tasks, even when the model is not explicitly trained for them.
@@
Our CT-RATE dataset, comprising paired chest CT volumes and radiology text reports, is publicly accessible via https://huggingface.co/datasets/ibrahimhamamci/CT-RATE [67]. The external validation set, RAD-ChestCT, is also openly accessible through https://zenodo.org/records/6406114 [68].
@@
Our trained models and codebase are publicly available at https://github.com/ibrahimethemhamamci/CT-CLIP [69] and https://github.com/ibrahimethemhamamci/CT-CHAT [70] for further research.
Acknowledgments
We extend our sincere gratitude to the Helmut Horten Foundation for their generous support, which made this work possible. We also thank Istanbul Medipol University for their invaluable support and provision of data. Some of the HPC resources were provided by the Erlangen National High Performance Computing Center (NHR@FAU) at Friedrich-Alexander-Universität Erlangen-Nürnberg under NHR projects b143dc and b180dc. NHR is funded by federal and Bavarian state authorities, and the NHR@FAU hardware is partly funded by the German Research Foundation (DFG; grant 440719683). This research was supported in part by the Intramural Research Program of the National Library of Medicine (NLM), National Institutes of Health (NIH), and used the computational resources of the NIH high-performance computing Biowulf cluster. B.K. received research support from the ERC project MIA-NORMAL 101083647, DFG grants 513220538 and 512819079, and the state of Bavaria (HTA). C.B. received research support from the Promedica Foundation, Chur, Switzerland. C.W., W.D., and K.B. are supported by NIH grant R01 HL141813, the Hariri Institute for Computing’s Junior Faculty Fellows Program, and its Focused Research Program. Figures 1, 3, 5, and 6 were created with BioRender.com.